# Must Read: A Systematic Survey of Computational Persuasion

### v1

@@ 7.1.3. Long Context Persuasion
(1) Benchmarking Long-Horizon Persuasion: Existing evaluation methods focus on short interactions, leaving a gap in assessing persuasive success over extended conversations. Future benchmarks should track how persuasive influence evolves across multiple turns and sessions, considering factors like memory, trust dynamics, and user resistance over time.
@@ 7.2. AI as Persuadee
While recent work has shown that LLMs can be susceptible to persuasive prompts and adversarial dialogue strategies (Zeng et al., 2024; Xu et al., 2024a; Bozdag et al., 2025), the boundaries and underlying mechanisms of these vulnerabilities are not well understood. Several important directions merit further investigation.
@@
(2) Effective Modalities of Agent Persuasion: While natural language is the dominant input modality, it is not yet known whether alternative channels such as structured metadata, tool outputs, or interleaved code could more effectively persuade models. Exploring multimodal input strategies may uncover novel vectors of influence and corresponding risks.
@@
(1) Defining Selectiveness in Persuasion: Persuasion resistance should not emerge from general stubbornness or degraded instruction-following, but from correctly identifying and rejecting problematic persuasive attempts. We must establish clear definitions and metrics for selective robustness that distinguish between appropriate and inappropriate influence.
@@ 7.3.1. Detecting & Evaluating Persuasion
(3) Predicting User Responses to Persuasion: Beyond identifying persuasive strategies, models should aim to predict how users are likely to respond to persuasive messages. Anticipating user reactions would enable more adaptive and context-sensitive evaluations, supporting downstream applications such as safer generation, persuasion mitigation, and user-personalized feedback.
@@
While models are already challenged in detecting single-turn persuasive attempts, the risks associated with long-term, contextually embedded persuasion are likely to grow. As conversational agents become more integrated into daily life, retain memory of user preferences, beliefs, and personality traits, they may pose subtle but serious risks by building trust over time and later leveraging it to influence users in ways that may not align with their best interests. Main challenges in this direction are:
@@ 8. CONCLUSION
In this survey of AI-driven persuasion, we have systematically examined existing research through the lens of AI as Persuader, AI as Persuadee, and AI as Persuasion Judge, with a focus on three core areas: Evaluating Persuasion, Generating Persuasion, and Safeguarding Persuasion. As persuasion becomes an increasingly inherent capability of large language models, research must prioritize developing robust methods for evaluating persuasive behavior, responsibly generating persuasive content, and carefully safeguarding against the potential risks posed by advanced AI systems. Our survey highlights that persuasion in AI systems poses challenges not only for human users but also within agentic ecosystems where models interact and influence one another. Looking ahead, we have outlined several promising future directions in this rapidly growing area of research, including improved and unified evaluation mechanisms, adaptive persuasive systems, and techniques for selective acceptance of persuasion. While our survey focuses primarily on English text-based persuasion, we acknowledge the importance of expanding this work to cover multilingual and multimodal contexts, which remain underexplored. We call on the research community to explore these diverse dimensions of persuasion in order to build safer, more ethical, and more trustworthy AI systems.

### v2

@@ 7.1.3. Long Context Persuasion
(1) Benchmarking Long-Horizon Persuasion: Existing evaluation methods focus on short interactions, not assessing persuasive success over extended conversations. Future benchmarks should track how persuasive influence evolves across multiple turns and sessions, considering factors like memory, trust dynamics, and user resistance over time.
@@ 7.2. AI as Persuadee
While recent work has shown that LLMs can be susceptible to persuasive prompts and adversarial strategies (Zeng et al., 2024; Xu et al., 2024a; Bozdag et al., 2025), the boundaries and underlying mechanisms of these vulnerabilities are not well understood. Insights from steering vectors may further clarify such mechanisms (Dunefsky and Cohan, 2025), and several important directions merit further investigation.
@@
(2) Effective Modalities of Agent Persuasion: While natural language is the dominant input modality, it is not yet known whether alternatives such as structured metadata, tool outputs, or interleaved code could more effectively persuade models. Exploring multimodal input may uncover novel vectors of influence and corresponding risks.
@@
(1) Defining Selectiveness in Persuasion: Persuasion resistance should not emerge from general stubbornness or degraded instruction-following, but from correctly identifying problematic persuasive attempts. We must establish clear definitions and metrics for selective robustness that distinguish between appropriate and inappropriate influence.
@@ 7.3.1. Detecting & Evaluating Persuasion
(3) Predicting User Responses to Persuasion: Models should aim to predict how users are likely to respond to persuasive messages. Anticipating user reactions would enable more adaptive and context-sensitive evaluations, supporting downstream applications such as safer generation, persuasion mitigation, and user-personalized feedback.
@@
Detecting persuasion across long-term interactions presents distinct challenges beyond single-turn analysis. Rather than simply recognizing persuasive intent in isolated exchanges, judge models must discern how influence evolves gradually through accumulated trust, personalization, and subtle framing. As conversational agents retain user-specific memories and adapt, identifying when a conversation shifts from rapport-building to strategic persuasion becomes increasingly complex. Main challenges in this direction are:
@@ 8. CONCLUSION
In this survey of AI-driven persuasion, we have systematically examined existing research through the lens of AI as Persuader, AI as Persuadee, and AI as Persuasion Judge, with a focus on three core areas: Evaluating Persuasion, Generating Persuasion, and Safeguarding Persuasion. As persuasion becomes an increasingly inherent capability of large language models, research must prioritize developing robust methods for evaluating persuasive behavior, responsibly generating persuasive content, and carefully safeguarding against the potential risks posed by advanced AI systems. Our survey highlights that persuasion in AI systems poses challenges not only for human users but also within agentic ecosystems where models interact and influence one another. Looking ahead, we have outlined several promising future directions in this rapidly growing area of research, including improved and unified evaluation mechanisms, adaptive persuasive systems, and techniques for selective acceptance of persuasion. While our survey focuses primarily on English text-based persuasion, we acknowledge the importance of expanding this work to cover multilingual and multimodal contexts, which remain underexplored. Moreover, a richer interdisciplinary integration with fields such as psycholinguistics, behavioral psychology, and communication studies can offer valuable insights into phenomena like argument framing, behavior change, and language processing, thereby informing the design, evaluation, and ethical grounding of persuasive AI systems. We call on the research community to explore these diverse dimensions of persuasion to build safer, more ethical, and more trustworthy AI systems.